If someone takes two tests and the highest grade was 71 points higher than the lowest grade. The sum of the two tests is 125,hat would be the lowest grade? Could you show work too so then I can get a clear view?
Answer by richwmiller(17219)

x+x+71=125
2x+71=125
2x=54
x=27
x+71=98